Provide a legend, equation, solution and conclusion for the following question:
Find three consecutive even integers whose sum is 114.

Answer by rfer(16322)

first x
second x+2
third x+4
----------------
x+x+2+x+4=114
3x=114-6
3x=108
x=36
x+2=38
x+4=40